How many committes can be formed from a group of 10 senators if each committe consists of 5 senators.
Answer by checkley77(12844)

10/5=2 COMMITTEES CAN BE FORMED IF MIX IS UNIMPORTANT.
IF YOU'RE LOOKING FOR SPECIFIC MIXES THEN:
(10*9*8*7*6)+1=30,240+1=30,241 DIFFERENT COMMITTEES CAN BE FORMED.